# VoiceShop: A Unified Speech-to-Speech Framework for Identity-Preserving Zero-Shot Voice Editing

### 10 Apr 2024 v1

@@ 3.3.2 Bottleneck-to-Bottleneck (BN2BN) Modeling for Many-to-Many Accent and Speech Style Conversion
We propose a bottleneck-to-bottleneck (BN2BN) model capable of many-to-many accent and speech style conversion by mapping the time-varying “bottleneck” content features of utterances from an arbitrary number of source accents to those of an arbitrary number of target accents in a single model using a multi-decoder architecture based on encoder-decoder sequence-to-sequence modeling with cross attention, effectively reducing the accent conversion task to a machine translation problem. Beyond typical accent conversion, the BN2BN model is also capable of generalized speech style transfer. There are no specific requirements for what constitutes a “speech style,” which may be as broad as emotional speech, the speaking styles of iconic personalities from popular culture, and so on.
@@
Our approach is rooted in the observation that these latent sequences not only encode the semantics of a speech signal (i.e., what is said), but also pronunciation and prosody (i.e., how it is said). During training, each batch $X\in\mathbb{R}^{B\times T\times D_{1}}$ consists of the local content features of utterances, such that the source accent, content, and speaker of inputs are sampled uniformly at random. The utterances are processed by a universal encoder $E$ that learns accent-agnostic latent representations $Z=E(X)\in\mathbb{R}^{B\times T\times D_{2}}$, which are passed as input to autoregressive domain-specific decoders $D_{t_{j}}$ and post nets $P_{t_{j}}$ for each supported target accent $t_{j}\in T$ to reconstruct the content features in the respective target accents regardless of the source accent, i.e., $P_{t_{j}}(D_{t_{j}}(Z))=\hat{X}_{t_{j}}\in\mathbb{R}^{B\times T\times D_{1}}$, as depicted in Figure 3.
@@
We adopt a conformer encoder and LSTM decoders with the same configuration proposed in Tacotron 2 (Shen et al., 2018), using additive energy-based Dynamic Convolution Attention (DCA) (Battenberg et al., 2020) as the cross-attention mechanism between the encoder and decoders. We jointly train “stop” gate projection layers for each decoder using binary CE, which produces a scalar between 0 and 1 indicating when to stop generation ($\mathcal{L}_{\text{Gate}}$). Based on the selection of time-varying content features, we apply three MAE reconstruction loss terms applied to the output of the LSTM decoders ($\mathcal{L}_{\text{Pre L10}}$), post nets ($\mathcal{L}_{\text{Post L10}}$), and universal predictor ($\mathcal{L}_{\text{L18}}$). The final loss is formulated as:
@@
We train each BN2BN model using the AdamW optimizer (Loshchilov & Hutter, 2019) on 8 A100 GPUs with a batch size of 32 samples and an initial learning rate of $1\times 10^{-4}$, to which we apply a decay rate of 0.85 every 100 epochs. The rate of convergence typically relies on the size of the dataset and varies between approximately 75 to 200 epochs.
@@
YourTTS (Casanova et al., 2022) 2.89$\pm$0.07 3.04$\pm$0.07
DiffVC (Popov et al., 2021) 3.14$\pm$0.07 3.42$\pm$0.07
@@
4 Experiments and Analysis
Evaluation Metrics. We evaluate the performance of each mentioned synthesis task using a variety of subjective and objective metrics, defined as follows.
To objectively evaluate the speaker similarity between ground truth input and synthesized output pairs, we use the Automatic Speaker Verification (ASV) metric, whereby we compute the cosine similarity of fixed-length embeddings extracted from a pre-trained speaker verification model. Therefore, the ASV metric falls within the range of -1 and 1, where greater values indicate higher similarity within the latent space of the speaker verification model, which is assumed to effectively discriminate timbre. We follow VALL-E (Wang et al., 2023a) and Voicebox (Le et al., 2023) and adopt WavLM-TDNN [^6] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification (Chen et al., 2022) to extract embeddings for this purpose. Additional task-specific metrics are defined in subsequent sections as needed.
For tasks where no obvious third-party baselines exist for direct comparison, such as style conversion and cross-lingual AC, we report scores of intentionally incorrect pairs (i.e., for ASV, we compute the similarity of non-matching speaker pairs). When placed alongside the scores of ground truth inputs, doing so establishes upper and lower bounds against which we can more meaningfully understand the values produced by VoiceShop’s outputs. For all applicable metrics, we summarize results with 95% confidence intervals.
4.1 Zero-Shot Voice Conversion
We demonstrate VoiceShop’s performance on monolingual and cross-lingual zero-shot VC, which refers to the task of modifying the timbre of a spoken utterance while preserving its content for arbitrary speakers not seen during training. In the monolingual case, both the source content and target timbre are spoken in the same language, whereas in the cross-lingual case, they are from different languages (e.g., apply the timbre of a Mandarin utterance to English content). Since the global speaker embedding used in our diffusion backbone model collapses temporal information, we additionally show that conversion can be achieved for out-of-domain languages not seen by the diffusion backbone model during training, enabling anyone to speak fluent English and Mandarin in their own voice regardless of their native language.
Subjective MOS Evaluation. We evaluate our framework on zero-shot VC against YourTTS (Casanova et al., 2022) and DiffVC (Popov et al., 2021), two recent works regarded as SOTA in this task. Following the experimental setups of these works, we use the VCTK corpus (Veaux et al., 2016) as the test set. We identify 8 VCTK speakers (4 males and 4 females) that are held out by both works, while our model is not trained on any VCTK data. We select one speech sample for each speaker as their reference, then each VC model will convert the voice of each speaker to the voice of all other speakers, yielding 56 samples per model. As ground truth for all test subsets, we randomly select 7 more audios for each of the test speakers, creating another 56 samples.
We conduct a MOS study with third-party vendors, recruiting 20 native English speakers (12 males, 8 females, mean age 32.0 years with a standard deviation of 6.5 years) for the study, in which each voice-converted sample or ground truth sample is presented together with the reference sample. We ask participants to rate the sample’s speech naturalness independently, then rate the sample’s speaker similarity with respect to the reference sample, disregarding its naturalness. We adopt a five-point Likert scale for both questions.
We summarize the MOS study results in Table 3. In both dimensions, our VoiceShop model outperforms other baseline models by clear margins.
Table 4: Results of monolingual and cross-lingual zero-shot voice conversion objective evaluation, measuring speaker similarity. We denote the number of individual samples used to calculate each value in parentheses.
Conversion Type ASV (normal-$\uparrow$)
English-to-English 0.435$\pm$0.005 (2,000)
Mandarin-to-Mandarin 0.756$\pm$0.002 (2,000)
English-to-Mandarin 0.340$\pm$0.003 (2,000)
Mandarin-to-English 0.407$\pm$0.003 (2,000)
OOD Language-to-English 0.334$\pm$0.004 (2,000)
OOD Language-to-Mandarin 0.389$\pm$0.006 (2,000)
Overall 0.438$\pm$0.003 (12,000)
Ground Truth 0.883$\pm$0.004 (600)
Non-Matching Speaker 0.127$\pm$0.004 (4,350)
Objective Evaluation. We use ASV to further evaluate the speaker similarity of our framework, as specified above. In particular, we generate input speech in English and Mandarin using Microsoft Azure TTS models in 10 timbres and synthesize an additional 10 utterances per target language to serve as the source content. We measure the quality of our model on monolingual zero-shot VC for English and Mandarin, cross-lingual zero-shot VC for English and Mandarin (both of which are in-domain to our model), and cross-lingual zero-shot VC for out-of-domain (OOD) languages applied to English and Mandarin content (we generate speech in Arabic, French, German, Greek, Hindi, Italian, Korean, Spanish, Tamil, and Zulu for this purpose). For these samples, we use our VS-EN-CN-L10 backbone model to generate speech and summarize our results in Table 4.
@@ 4.2 Zero-Shot Identity-Preserving Many-to-Many Accent Conversion
Subjective CMOS Evaluation. We evaluate our model against Jin et al. (2023), a recent work which also achieves zero-shot identity-preserving many-to-many accent conversion, curating 16kHz samples directly from their official demo page [^9] https://accent-conversion.github.io/. As our diffusion backbone model requires 24kHz input, we use the official pre-trained checkpoint of AudioSR-Speech [^10] https://github.com/haoheliu/versatile_audio_super_resolution, a versatile audio super-resolution model, to predict the corresponding high-fidelity 48kHz speech, which we then downsample to 24kHz prior to performing inference with our framework. Since upsampling the 16kHz audio is lossy and cannot recover spectral information between 8 to 12kHz in the frequency domain, it is necessary to use a generative model to properly extend the bandwidth. After inference, all samples are downsampled to 16kHz for a fair comparison.
We conduct two CMOS studies to respectively measure the accent strength (aCMOS) and speaker similarity (sCMOS) of our systems using 30 pairs of samples (specifically, we evaluate on their British-to-American, British-to-Indian, and Indian-to-American subsets). Both studies recruit 20 anonymous native English speakers balanced across gender using the Prolific crowd-sourcing platform [^11] https://www.prolific.com/, where participants are asked to rate pairs of samples on a scale from -3 to 3, where -3 indicates a strong preference for the baseline model, 3 indicates a strong preference for VoiceShop, and 0 indicates no preference. Each comparison is made against provided reference clips to standardize the comparisons between listeners, who may have different expectations of the target accents’ characteristics. The results are summarized in Table 5. We find that participants prefer the accent strength of our system by a clear margin, while maintaining relatively equal performance on speaker similarity.
@@
Objective Evaluation. We conduct additional objective evaluations to judge VoiceShop’s monolingual and cross-lingual AC capabilities. Specifically, we train three accent classifiers consisting of an ECAPA-TDNN encoder (Desplanques et al., 2020) and multilayer perceptron on the same datasets used to train our BN2BN models for monolingual English AC, monolingual Mandarin AC, and cross-lingual AC. All classifiers are trained using AdamW (Loshchilov & Hutter, 2019) for 24H on 8 V100 GPUs with a batch size of 64, a learning rate of $1\times 10^{-4}$, and decay rate of 0.85 applied every 100 epochs, achieving validation accuracies of 99.3% in English, 99.8% in Mandarin, and 99.9% in the cross-lingual case on their respective holdout sets.
For each conversion type, we synthesize input speech in a variety of accents using the Microsoft Azure TTS service and convert each source accent to each target accent using our models. We use the learned latent space of each classifier to investigate the accent similarity of samples generated by our system, first computing the centroids of embeddings extracted from the last layer of each classifier using reference samples for each target accent. We then extract embeddings of samples to calculate the cosine similarity against the averaged reference embeddings for three categories: a separate set of ground truth samples, the accent-converted outputs of our models, and non-matching accents to establish a performance lower bound. We report overall similarities for each conversion type in Table 6 and provide a detailed breakdown of accent-wise similarities in Table A5.
@@
Table 7: Results of monolingual and cross-lingual accent conversion objective evaluation, measuring speaker similarity. We denote the number of individual samples used to calculate each value in parentheses.
Conversion Type ASV (normal-$\uparrow$)
English Monolingual 0.508$\pm$0.006 (2,000)
Mandarin Monolingual 0.809$\pm$0.003 (1,280)
Cross-Lingual 0.625$\pm$0.010 (1,280)
Overall 0.625$\pm$0.005 (4,560)
Ground Truth 0.925$\pm$0.001 (1,900)
Non-Matching Speaker 0.175$\pm$0.005 (3,420)
To visualize these results, we generate t-SNE plots (van der Maaten & Hinton, 2008) of the predicted embeddings of input speech in each source accent and output speech in each target accent, as depicted in Figure 5. We find the accent classifiers effectively cluster input speech according to their source accents and further observe that output speech generally preserves these clusters after accent conversion regardless of the source accent.
@@
Figure 5: By using the latent space of accent classifiers, we observe that accent-converted speech predicted by our BN2BN models produce clusters according to their target accents.
We evaluate the speaker similarity of our BN2BN method using ASV and summarize these results in Table 7.
@@ 4.3 Zero-Shot Identity-Preserving Speech Style Conversion
We evaluate VoiceShop’s speech style conversion capabilities using three styles (“Sarcastic Youth,” “Formal British,” and “Cartoon Character”). In these examples, the target styles are acquired from in-house datasets, whereby actors perform highly stylized speech, amounting to a total of approximately one hour of data per style. Conversion occurs in a many-to-one manner, such that English speakers of arbitrary source accents are converted to the target style while preserving their timbre in a zero-shot manner. We use our VS-EN-L10 backbone model to generate samples. To our knowledge, VoiceShop is the first framework to extend AC to the generalized speech style conversion task.
Subjective MOS Evaluation. To measure the perceived conversion strength of the aforementioned speech styles, we conduct a MOS study consisting of 20 anonymous native English speakers balanced across gender using the Prolific crowd-sourcing platform. We use the Microsoft Azure TTS service to generate speech in 7 timbres and convert each to the target speaking styles using our BN2BN models. We ask participants to rate the perceived similarity of a sample’s speaking style against a provided reference clip on a scale from 1 to 5, where 1 indicates low similarity and 5 indicates high similarity, for three categories: ground truth samples extracted from the training corpus, neutral speech inputs, and model outputs with converted speaking styles (21 samples each). We summarize results in Table 8 and find the outputs of our style conversion models perform comparably well to ground truth samples, as opposed to the neutral inputs, indicating that the target speaking styles are reliably transferred to unseen speakers.
Table 8: Results of subjective MOS evaluation of style conversion measuring conversion strength of three styles (“Sarcastic Youth,” “Formal British,” and “Cartoon Character”) compared against ground truth samples.
Model MOS (normal-$\uparrow$)
Model Input 1.57$\pm$0.10
@@
Ground Truth 4.05$\pm$0.12
Objective Evaluation. We again use ASV to measure the speaker similarity of our style conversion models under the same configuration and provide results in Table 9.
@@
Conversion Type ASV (normal-$\uparrow$)
Ground Truth 0.903$\pm$0.003 (210)
Model Output 0.492$\pm$0.015 (210)
Non-Matching Speaker 0.168$\pm$0.012 (420)

### 11 Apr 2024 v2

@@ 3.3.2 Bottleneck-to-Bottleneck (BN2BN) Modeling for Many-to-Many Accent and Speech Style Conversion
We propose a bottleneck-to-bottleneck (BN2BN) model capable of many-to-many accent and speech style conversion by mapping the time-varying “bottleneck” content features of utterances from an arbitrary number of source accents to those of an arbitrary number of target accents in a single model using a multi-decoder architecture based on encoder-decoder sequence-to-sequence modeling with cross attention, effectively reducing the accent conversion task to a machine translation problem. Beyond typical accent conversion, the BN2BN model is also capable of generalized speech style transfer. There are no specific requirements for what constitutes a “speech style,” which may be as broad as emotional speech or the speaking styles of iconic personalities from popular culture.
@@
Our approach is rooted in the observation that these latent sequences not only encode the semantics of a speech signal (i.e., what is said), but also pronunciation and prosody (i.e., how it is said). During training, each batch $X\in\mathbb{R}^{B\times T\times D_{1}}$ consists of the local content features of utterances, such that the source accent, content, and speaker of inputs are sampled uniformly at random. The utterances are processed by a universal encoder $E$ that learns accent-agnostic latent representations $Z=E(X)\in\mathbb{R}^{B\times T\times D_{2}}$, which are passed as input to autoregressive domain-specific decoders $D_{j}$ and post nets $P_{j}$ for each supported target accent $j\in[1,M]$ to reconstruct the content features in the respective target accents regardless of the source accent, i.e., $P_{j}(D_{j}(Z))=\hat{X}_{j}\in\mathbb{R}^{B\times T\times D_{1}}$, as depicted in Figure 3.
@@
We adopt a conformer encoder and LSTM decoders with the same configuration proposed in Tacotron 2 (Shen et al., 2018), using additive energy-based Dynamic Convolution Attention (DCA) (Battenberg et al., 2020) as the cross-attention mechanism between the encoder and decoders. We jointly train “stop” gate projection layers for each decoder using binary CE, which produces a scalar between 0 and 1 indicating when to stop generation ($\mathcal{L}_{\text{Gate}}$). We apply three MAE reconstruction loss terms to the output of the LSTM decoders ($\mathcal{L}_{\text{Pre L10}}$), post nets ($\mathcal{L}_{\text{Post L10}}$), and universal predictor ($\mathcal{L}_{\text{L18}}$). The final loss is formulated as:
@@
We train each BN2BN model using the AdamW optimizer (Loshchilov and Hutter, 2019) on 8 A100 GPUs with a batch size of 32 samples and an initial learning rate of $1\times 10^{-4}$, to which we apply a decay rate of 0.85 every 100 epochs. The rate of convergence typically relies on the size of the dataset and varies between approximately 75 to 200 epochs.
4 Experiments and Analysis
Evaluation Metrics.
We evaluate the performance of each mentioned synthesis task using a variety of subjective and objective metrics, defined as follows.
To objectively evaluate the speaker similarity between ground truth input and synthesized output pairs at a large scale, we use the Automatic Speaker Verification (ASV) metric, whereby we compute the cosine similarity of fixed-length embeddings extracted from a pre-trained speaker verification model. Therefore, the ASV metric falls within the range of -1 and 1, where greater values indicate higher similarity within the learned latent space of the speaker verification model, which is assumed to effectively discriminate timbre. We follow VALL-E (Wang et al., 2023a) and Voicebox (Le et al., 2023) and adopt WavLM-TDNN [^6] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification (Chen et al., 2022) to extract speaker embeddings for this purpose. Additional task-specific metrics are defined in subsequent sections as needed.
For tasks where no obvious third-party baselines exist for direct comparisons, such as style conversion and cross-lingual AC, we report scores of intentionally incorrect pairs (i.e., for ASV, we compute the similarity of non-matching speaker pairs). When placed alongside the scores of ground truth inputs, doing so establishes upper and lower bounds against which we can more meaningfully understand the values produced by VoiceShop’s outputs. For all applicable metrics, we summarize results with 95% confidence intervals.
4.1 Zero-Shot Voice Conversion
We demonstrate VoiceShop’s performance on monolingual and cross-lingual zero-shot VC, which refers to the task of modifying the timbre of a spoken utterance while preserving its content for arbitrary speakers not seen during training. In the monolingual case, both the source content and target timbre are spoken in the same language, whereas in the cross-lingual case, they are from different languages (e.g., apply the timbre of a Mandarin utterance to English content). Since the global speaker embedding used in our diffusion backbone model collapses temporal information, we additionally find that conversion can be achieved for out-of-domain languages not seen by the diffusion backbone model during training, enabling anyone to speak fluent English and Mandarin in their own voice regardless of their native language.
Subjective MOS Evaluation.
We evaluate our framework on zero-shot VC against YourTTS (Casanova et al., 2022) and DiffVC (Popov et al., 2021), two recent works regarded as SOTA in this task. Following the experimental setups of these works, we use the VCTK corpus (Veaux et al., 2016) as the test set. We identify 8 VCTK speakers (4 males and 4 females) that are held out by both works, while our model is not trained on any VCTK data. We select one speech sample for each speaker as their reference, then each VC model will convert the voice of each speaker to the voice of all other speakers, yielding 56 samples per model. As ground truth for all test subsets, we randomly select 7 more audios for each of the test speakers, creating another 56 samples.
@@ 4.1 Zero-Shot Voice Conversion
YourTTS 2.89$\pm$0.07 3.04$\pm$0.07
DiffVC 3.14$\pm$0.07 3.42$\pm$0.07
@@
We conduct a MOS study with third-party vendors, recruiting 20 native English speakers (12 males, 8 females, mean age 32.0 years with a standard deviation of 6.5 years) for the study, in which each voice-converted sample or ground truth sample is presented together with the reference sample. We ask participants to rate the sample’s speech naturalness independently, then rate the sample’s speaker similarity with respect to the reference sample, disregarding its naturalness. We adopt a five-point Likert scale for both questions. We summarize the MOS study results in Table 3. In both dimensions, our VoiceShop model outperforms other baseline models by clear margins.
@@ 4.2 Zero-Shot Identity-Preserving Many-to-Many Accent Conversion
Subjective CMOS Evaluation.
We evaluate our model against Jin et al. (2023), a recent work which also achieves zero-shot identity-preserving many-to-many accent conversion, curating 16kHz samples directly from their official demo page [^9] https://accent-conversion.github.io/. As our diffusion backbone model requires 24kHz input, we use the official pre-trained checkpoint of AudioSR-Speech [^10] https://github.com/haoheliu/versatile_audio_super_resolution, a versatile audio super-resolution model, to predict the corresponding high-fidelity 48kHz speech, which we then downsample to 24kHz prior to performing inference with our framework. Since upsampling the 16kHz audio is lossy and cannot recover spectral information between 8 to 12kHz in the frequency domain, it is necessary to use a generative model to properly extend the bandwidth. After inference, all samples are downsampled to 16kHz for a fair comparison.
@@
We conduct two CMOS studies to respectively measure the accent strength (aCMOS) and speaker similarity (sCMOS) of our systems using 30 pairs of samples (specifically, we evaluate on their British-to-American, British-to-Indian, and Indian-to-American subsets). Both studies recruit 20 anonymous native English speakers balanced across gender using the Prolific crowd-sourcing platform [^11] https://www.prolific.com/, where participants are asked to rate pairs of samples on a scale from -3 to 3, where -3 indicates a strong preference for the baseline model, 3 indicates a strong preference for VoiceShop, and 0 indicates no preference. Each comparison is made against provided reference clips to standardize the comparisons between listeners, who may have different expectations of the target accents’ characteristics. The results are summarized in Table 4. We find that participants prefer the accent strength of our system by a clear margin while maintaining relatively equal performance on speaker similarity.
Objective Evaluation.
We conduct additional objective evaluations to judge VoiceShop’s monolingual and cross-lingual AC capabilities. Specifically, we train three accent classifiers consisting of an ECAPA-TDNN encoder (Desplanques et al., 2020) and multilayer perceptron on the same datasets used to train our BN2BN models for monolingual English AC, monolingual Mandarin AC, and cross-lingual AC. All classifiers are trained using AdamW (Loshchilov and Hutter, 2019) for 24H on 8 V100 GPUs with a batch size of 64, a learning rate of $1\times 10^{-4}$, and decay rate of 0.85 applied every 100 epochs, achieving validation accuracies of 99.3% in English, 99.8% in Mandarin, and 99.9% in the cross-lingual case on their respective holdout sets.
Table 5: Results of monolingual and cross-lingual accent conversion objective evaluation, measuring speaker similarity. We denote the number of samples used to calculate each value in parentheses.
Conversion Type ASV (normal-$\uparrow$)
English Monolingual 0.508$\pm$0.006 (2,000)
Mandarin Monolingual 0.809$\pm$0.003 (1,280)
Cross-Lingual 0.625$\pm$0.010 (1,280)
Overall 0.625$\pm$0.005 (4,560)
Same Speaker 0.925$\pm$0.001 (1,900)
Non-Matching Speaker 0.175$\pm$0.005 (3,420)
For each conversion type, we synthesize input speech in a variety of accents using the Microsoft Azure TTS service and convert each source accent to each target accent using our models. We use the learned latent space of each classifier to investigate the accent similarity of samples generated by our system, first computing the centroids of embeddings extracted from the last layer of each classifier using reference samples for each target accent. We then extract embeddings of samples to calculate the cosine similarity against the averaged reference embeddings for three categories: a separate set of ground truth samples, the accent-converted outputs of our models, and non-matching accents to establish a performance lower bound. We report overall similarities in Table 6 and provide a detailed breakdown of accent-wise similarities in Table A5.
@@
To visualize these results, we generate t-SNE plots (van der Maaten and Hinton, 2008) of the predicted embeddings of input speech in each source accent and output speech in each target accent, as depicted in Figure 5. We find the accent classifiers effectively cluster input speech according to their source accents and further observe that output speech generally preserves these clusters after accent conversion regardless of the source accent.
We evaluate the speaker similarity for each conversion type of our BN2BN method alongside additional reference values using the ASV metric. We find that the timbre of the original speaker is largely preserved in both monolingual and cross-lingual scenarios after accent conversion, as summarized in Table 5.
@@
Figure 5: Visualizing accent transfer: By using the latent space of accent classifiers, we observe that input speech is clustered by source accent and that accent-converted speech predicted by our BN2BN models largely preserve these structures according to their target accents.
@@ 4.3 Zero-Shot Identity-Preserving Speech Style Conversion
We evaluate VoiceShop’s speech style conversion capabilities using three styles (“Sarcastic Youth,” “Formal British,” and “Cartoon Character”). In these examples, the target styles are acquired from in-house datasets, whereby actors perform highly stylized speech, amounting to a total of approximately one hour of data per style. Conversion occurs in a many-to-one manner, such that English speakers of arbitrary source accents are converted to the target style while preserving their timbre in a zero-shot manner. We use our VS-EN-L10 backbone model to generate samples.
Subjective MOS Evaluation.
To measure the perceived conversion strength of the aforementioned speech styles, we conduct a MOS study consisting of 20 anonymous native English speakers balanced across gender using the Prolific crowd-sourcing platform. We use the Microsoft Azure TTS service to generate speech in 7 timbres and convert each to the target speaking styles using our BN2BN models. We ask participants to rate the perceived similarity of a sample’s speaking style against a provided reference clip on a scale from 1 to 5, where 1 indicates low similarity and 5 indicates high similarity, for three categories: ground truth samples extracted from the training corpus, neutral speech inputs, and model outputs with converted speaking styles (21 samples each). We summarize results in Table 8 and find the outputs of our style conversion models perform comparably well to ground truth samples, as opposed to the neutral inputs, indicating that the target speaking styles are reliably transferred to unseen speakers.
Table 7: Results of subjective MOS evaluation of style conversion measuring conversion strength of three styles (“Sarcastic Youth,” “Formal British,” and “Cartoon Character”).
Speech Type Style MOS (normal-$\uparrow$)
Non-Stylized Input 1.57$\pm$0.10
@@
Same Speaker and Style 4.05$\pm$0.12
Speech Type ASV (normal-$\uparrow$)
Non-Matching Speaker 0.168$\pm$0.012 (420)
Model Output 0.492$\pm$0.015 (210)
Same Speaker and Style 0.903$\pm$0.003 (210)
@@
Objective Evaluation.
We again use ASV to measure the speaker similarity of our style conversion models under the same configuration and provide results in Table 8. We note that a fair comparison should be made by calculating the ASV between the same speaker’s stylized and non-stylized speech. However, such paired samples do not exist as the ground truth. Therefore, the ASV score for same speaker and style in Table 8 does not expose the ASV drop due to the stylization of the same person’s speech, which is reflected in the score for our model output.